• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dohrmann

    Log entry time 00:00:07 on February24,2001

    Entry number 56888

    keyword=Shift summary

    Took data in kinematic 15 , 4He target, radiator in and out.
    For run 2168 we set the pedestals to zero. For run 2169 we restored the
    pedestals.
    Beam was fine exect for a 45 min loss around 18:00

    FD
    runno Target Radiator Duration #Events Comments
    2158 4He out 30 min 146 K
    2159 Dummy out 30 min 357 K
    2160 Dummy in 25 min 1M with 47 uA beam


    2161 4He in 45 min 1M
    2162 4He in 45 min 1M

    2163 4He out 30 min 860K

    2164 4He in 50 min 1M
    2165 4He in 40 min 1M

    2166 4He out junk coda / roc reboot
    2167 4He out 30 min 930 K

    2168 4He in 60 min 800K Pedestals set to zero for this run 2169 4He in